# Junxiong Lin4, Jiahong Yan2, Jiawei Liu1, Zheng-Jun Zha1

### v2

@@ Appendix E Additional Experiments
E.1 Audio Transcript Length Analysis
As shown in Figure 14, we further analyze the impact of different audio transcript lengths (measured by the number of words) on model performance. A smaller number of audio words increases the risk of missing key information, while longer transcripts may introduce irrelevant noise, potentially confusing the reasoning process. We observe that using 50 audio words yields the best performance, suggesting that a moderate transcript length provides an effective balance between retaining essential information and avoiding unnecessary distraction.
E.2 GRPO Training Curves
@@
E.3 Consistency Analysis Between Judge Model and Human Evaluation
@@ E.4 Case Study
Appendix F Limitations
Finally, over-reliance on automated fact-checking systems carries significant societal risks. Misclassifying authentic content as misinformation can suppress legitimate discourse and harm the reputations or economic viability of content creators. Conversely, undetected misinformation—especially when reinforced by flawed but confident reasoning—may erode public trust, amplify polarization, and destabilize information ecosystems.
Appendix G Societal Impacts
Positive Societal Impacts
Negative Societal Impacts
Appendix H Responsible Release and Safeguards
@@
Appendix I Ethics Statement
This work includes human evaluation through crowdsourced annotation. Annotators were provided with clear task instructions and example interfaces. All participants gave informed consent, and their responses were anonymized. Compensation was provided in accordance with local fair-pay standards, and the study followed institutional ethical guidelines for human subjects research. We support transparency and reproducibility, and have made efforts to share data and experimental details responsibly, while prioritizing ethical considerations, participant privacy, and societal safety.
This study involved minimal risk to participants, limited to standard crowdsourced annotation tasks. All participants were informed of the nature and purpose of the task, their right to withdraw at any time, and the anonymous handling of their responses. No personally identifiable information (PII) was collected. Based on the low-risk nature of the study and in accordance with the guidelines of our institution, formal IRB approval was not required. The study nonetheless followed standard ethical principles for research involving human subjects.

### v3

@@ Appendix E Additional Experiments
E.1 GRPO Training Curves
@@
E.2 Consistency Analysis Between Judge Model and Human Evaluation
@@ E.3 Case Study
Appendix F Responsible Release and Safeguards
@@
Appendix G Ethics Statement
This work includes human evaluation via crowdsourced annotation. Annotators received clear task instructions and example interfaces, provided informed consent, and could withdraw at any time. Responses were anonymized, no personally identifiable information was collected, and compensation met or exceeded local fair-pay standards. The study adhered to institutional ethical guidelines and local labor regulations.
We collected a small annotation set (5k preference samples) under a transparent, compliant protocol. Given the non-sensitive content and minimal risk typical of standard annotation tasks, the study qualified for an IRB exemption under our institution’s policy. We support transparency and reproducibility and have shared data and experimental details responsibly, prioritizing ethical considerations, participant privacy, and societal safety.